# Unstable Unlearning: The Hidden Risk of Concept Resurgence in Diffusion Models

### v2

@@ 4.2 The Impact of Algorithmic Choices on Resurgence
The four algorithms we consider perform unlearning through fine-tuning model parameters, closed-form edits, or a combination of both. Fine-tuning optimizes an unlearning objective via gradient-based methods, as seen in ESD, which adjusts the model so that the score function conditioned on a concept matches the unconditional score function. Closed-form edits derive an explicit update for unlearning, as in UCE, which modifies key and value weights in cross-attention layers to replace concept-specific representations with generic or blank ones. MACE combines both approaches: it uses a closed-form edit to adjust word embeddings in concept-containing prompts and LoRA fine-tuning to suppress concept-related attention in generated images. We categorize ESD and SDD as fine-tuning methods, UCE as closed-form, and MACE as a hybrid approach.
@@
Finetuning vs. Closed-Form In Table 1, we see a gap in the severity of concept resurgence between the fine-tuning algorithms and those using closed-form edits. Specifically, UCE is quite robust, exhibiting very small resurgence across tasks. We conjecture that UCE is the strongest type of closed-form edit, as it modifies the cross attention weights to directly map the target concept to a higher-level (more abstract) concept. For example, if the target concept is a particular celebrity, it may be mapped to the more abstract concept like “a Person" or “a Celebrity". In contrast, MACE modifies the cross-attention weights to map the embeddings of all the surrounding words in the given prompts to be similar to embeddings of the surrounding words after replacing the target concept with a more abstract one. This difference means that MACE does not directly optimize the parameter update to move the target concept embedding towards the abstract concept embedding. Furthermore, because MACE incorporates unlearning the target concept information via fine-tuning, this might leave it more vulnerable to concept resurgence than UCE, which is based on a direct closed-form edit.
(a) Parameter choice in ESD algorithm to erase a single copyrighted character
(b) Scaling the ESD algorithm to erase a single celebrity
@@
Parameter Choice The second algorithmic factor we examine is which subsets of parameters are updated in the unlearning phase, and which (potentially overlapping) subsets of parameters are further fine-tuned. We start by showing how these choices potentially explain why UCE is more robust to concept resurgence than the other three algorithms. As discussed above, UCE only modifies the cross-attention weights with a closed form edit. As discussed in [14], this approach is very effective for concepts that are localized to the words themselves (e.g. the name of a celebrity; contrast this to unsafe content, which is a more abstract concept). Applying LoRA fine-tuning after UCE unlearning, we find no evidence of concept resurgence. We then instead fine-tune the full set of parameters, which yields a small degree of resurgence. Finally, motivated by this result, we choose to fully fine-tune the cross-attention layers only. We see that the resurgence is comparable between the two (Table 4), suggesting that the nature of UCE’s closed-form edit being localized to the cross-attention layers may make it very robust.
@@
The second difference between the four algorithms is the subset of model parameters that are updated in the unlearning step. Section 3 focuses primarily on modifying the cross-attention layers (with the exception of MACE, which also updates the rest of the model parameters via LoRA fine-tuning). Here, we focus on ESD in the single celebrity erasure task and the copyright erasure task, which both exhibit very high degrees of concept resurgence. In each of these tasks, we vary vary the subset of parameters that are updated in the unlearning step: either all of the parameters, all of the parameters except those in the cross-attention layers, and only those in the cross-attention layers. We find that the cross-attention parameters do indeed play the most important role in unlearning for these tasks and that unlearning on all the parameters only provided marginal gains in preventing resurgence (Fig. 4).
@@ 5 Why Does Concept Resurgence Occur?
Finally, to better understand the root cause(s) of concept resurgence, we explore this phenomenon in a simplified one-dimensional setting, where the distribution of interest is a simple mixture of standard Gaussians. This (intentionally stylized) model provides valuable intuition and insight into the dynamics of concept resurgence.
Setup. We construct two Gaussian distributions $p(x)$ and $q(x)$ with means $\mu$ and $\mu_{\text{FT}}$ respectively. We fix $\sigma^{2}=1$ for both distributions. $p(x)$ will model the original “pretraining" distribution, and $q(x)$ will model the distribution on which the model is fine-tuned. Next, we define a “concept" as the following membership function $c(x)=\mathbbm{1}[a\leq x\leq b]=1$ (i.e. a concept is represented as an interval on the real line). We model exact unlearning as (re)training a diffusion model on the data sampled from the original distribution, excluding values from the interval $c(x)$. To construct this distribution, we simply perform rejection sampling from $p(x)$, rejecting any samples which fall in the interval $c(x)$.
In this setting, we model an “approximate unlearning" algorithm as one which approximates the desired data distribution but leaves a probability mass of $\rho\in[0,1]$ on the unwanted interval. $\rho=0.0$ indicates exact unlearning and $\rho=1.0$ indicates no unlearning. We model approximate unlearning by simply training on a sample of data from $p(x)$ where we performed modified rejection sampling with a tolerance parameter of $\rho$ — if a sample lies in $c(x)$, it is rejected with probability $\rho$; otherwise, it is retained with probability $1$.
This setup allows us to investigate how varying levels of probability mass which remain in the unlearned concept region — corresponding to varying degrees of “success" in the initial unlearning step — can lead to concept resurgence. For the sake of this example, we start with a baseline level of approximate unlearning quality at $\rho\leq 30\%$. Finally, after applying unlearning, we fine-tune the resulting model on data sampled from $q(x)$ (after first rejecting any samples which lie in $c(x)$).
Figure 5: Impact of fine-tuning on concept resurgence in a one-dimensional setting. The concept to be unlearned is modeled as the interval $[-2,-1]$. The first plot depicts the true data distribution, excluding the unwanted concept. The second plot is the distribution learned by the diffusion model via exact unlearning. The third plot is the distribution learned by fine-tuning the model learned via exact unlearning. The non-zero probability left behind by exact unlearning on the unlearned concept is amplified by finetuning.
Training. With this setup, we train denoising score matching models (the same techniques used in the Stable Diffusion models studied previously) to model these distributions. Our diffusion models are based on the variance exploding SDE, where we choose the diffusion coefficient to be $g(t)=\lambda^{t}$. We train separate diffusion models for each value of $\rho$ to represent varying unlearning quality. Afterwards, we fine-tune each of these models on samples from $q(x)$. We use a KL divergence penalty in the score denoising loss when fine-tuning to prevent catastrophic forgetting.
The experiments we present are for the following setup: $p(x)\sim\mathcal{N}(-2.0,1.0)$, $q(x)\sim\mathcal{N}(-1.0,1.0)$ and $c(x)=\mathbbm{1}[-2.0\leq x\leq-1.0]$.
Figure 6: The average log-likelihood of five equally spaced values in the unlearned concept interval as unlearning quality increases (left). The number of samples (out of 10000 total) from the learned distribution that contain the unlearned concept. Fine-tuning introduces a constant resurgence.
Evaluation. To measure resurgence in this setting, we measure the average log-likelihood of five equally spaced points in the unwanted concept interval $[-2,-1]$ and the number of samples generated by the diffusion model that contain the concept. In practice, it is intractable to compute the log-likelihood of data under the learned distribution. However, because our model is one-dimensional, we can approximately compute the log-likelihoods via numerical integration. We provide more details about this procedure in Appendix G.
Results. We first consider exact unlearning, i.e. $\rho=0.0$. We plot the distributions of the original samples from both $p(x)$ and $q(x)$ with the learned distributions after (1) exact unlearning and (2) fine-tuning after exact unlearning in Figure 5. It is important to note that even under exact unlearning, the diffusion model leaves some non-zero probability mass on the unlearned concept region. We conjecture this is due to the implicit bias of diffusion models for learning smooth distributions (as also characterized by other works [1]), which leads to some mass being placed on the unwanted concept interval even though this region is outside the support of the training data. We further observe that fine-tuning amplifies this small amount of additional probability mass on the unlearned concept interval.
We now examine how this phenomenon changes as a function of the amount of probability mass which remains in the unwanted concept interval after unlearning. The degree of resurgence (as measured by the number of samples that contain the concept) is constant as we increase this probability (Figure 6). The average log likelihood also increases after fine-tuning, suggesting that one cause for resurgence might be the model’s inductive bias towards learning smooth distributions, which in turn places some small probability mass on the unwanted concept interval.
Although this mass may be negligible — so small that it is difficult to detect with sampling-based approaches, subsequent fine-tuning can lead to significant concept resurgence. This model is consistent with our empirical results, as well as those which appear elsewhere in the literature [13, 22, 13, 27] — unlearning algorithms typically suppress (rather than fully remove) the probability of generated an unwanted concept. Of course, this stylized model does not capture the full complexity of modern text-to-image diffusion models like Stable Diffusion v1.4 and v2.1, but our results shed light on possible factors driving concept resurgence and suggest avenues for future work.
@@ 6 Discussion and Limitations
The scale of generative models introduces new challenges, including the risk of learning concepts that are unsuitable or undesirable for certain downstream applications.
Ideally, unlearning algorithms would enable the precise and permanent removal of unwanted concepts while preserving the model’s overall utility. Reality, however, is more complex.
Our work uncovers a critical limitation of current unlearning methods, which we term concept resurgence. We demonstrate this phenomenon through rigorous empirical evaluations, highlighting the practical limitations of state-of-the-art unlearning techniques. These findings emphasize the need to rethink current approaches to concept erasure, especially in contexts where maintaining the integrity of model updates is essential.
Our investigation opens up several important avenues for future work. For example, we do not provide a theoretical characterization of concept resurgence, nor do we present a strategy designed to prevent it from happening. Both developments could help to enhance the robustness of unlearning methods. Additionally, our evaluations focus on a mix of well-known academic benchmarks and synthetic tasks, and further research is necessary to assess the prevalence of concept resurgence in real-world deployments (particularly the effect of interleaving a large number of compositional updates, which may exacerbate these vulnerabilities).
Concept resurgence also raises important questions about responsibility for downstream harms. Despite a developer’s best efforts to sanitize a model using these techniques, a downstream user who fine-tunes a published model might be surprised to discover that guardrails put in place by the developer no longer exist. This creates a dilemma: is the developer obligated to permanently and irrevocably erase problematic concepts, or does responsibility shift to the downstream if they (inadvertently) reintroduce them?
Despite these challenges, concept unlearning remains a valuable tool for model developers. By identifying its vulnerabilities, our work aims to drive the development of erasure techniques that remain robust throughout a model’s life-cycle or develop tools that can help developers anticipate when concept resurgence is likely to happen. Addressing these weaknesses will be essential for ensuring the safety and alignment of generative models as they are fine-tuned and adapted for diverse applications.
@@ Appendix A Additional Qualitative Examples
In this section we include qualitative results for the copyright and object erasure tasks in Figure 7 and Figure 8, respectively. These results are analogous to those presented in Figure 2 for the celebrity erasure task. We choose to exclude qualitative examples of resurgence for the unsafe content task, as these may be upsetting. For a quantitative evaluation of this task across all four unlearning algorithms, we refer readers to Table 1.

### v3

@@ 4.2 The Impact of Algorithmic Choices on Resurgence
The seven algorithms we consider perform unlearning through fine-tuning model parameters, closed-form edits, or a combination of both. Fine-tuning optimizes an unlearning objective via gradient-based methods, as seen in ESD, which adjusts the model so that the score function conditioned on a concept matches the unconditional score function. Closed-form edits derive an explicit update for unlearning, as in UCE, which modifies key and value weights in cross-attention layers to replace concept-specific representations with generic or blank ones. MACE combines both approaches: it uses a closed-form edit to adjust word embeddings in concept-containing prompts and LoRA fine-tuning to suppress concept-related attention in generated images. We categorize ESD and SDD as fine-tuning methods, UCE as closed-form, and MACE as a hybrid approach.
@@
Finetuning vs. Closed-Form In Table˜1, we see a gap in the severity of concept resurgence between the fine-tuning algorithms and those using closed-form edits. Specifically, UCE is quite robust, exhibiting very small resurgence across tasks. We conjecture that UCE is the strongest type of closed-form edit, as it modifies the cross attention weights to directly map the target concept to a higher-level (more abstract) concept. For example, if the target concept is a particular celebrity, it may be mapped to the more abstract concept like “a Person" or “a Celebrity". In contrast, MACE modifies the cross-attention weights to map the embeddings of all the surrounding words in the given prompts to be similar to embeddings of the surrounding words after replacing the target concept with a more abstract one. This difference means that MACE does not directly optimize the parameter update to move the target concept embedding towards the abstract concept embedding. Furthermore, because MACE incorporates unlearning the target concept information via fine-tuning, this might leave it more vulnerable to concept resurgence than UCE, which is based on a direct closed-form edit.
@@
Parameter Choice The second algorithmic factor we examine is which subsets of parameters are updated in the unlearning phase, and which (potentially overlapping) subsets of parameters are further fine-tuned. We start by showing how these choices potentially explain why UCE is more robust to concept resurgence than the other three algorithms. As discussed above, UCE only modifies the cross-attention weights with a closed form edit. As discussed in [14], this approach is very effective for concepts that are localized to the words themselves (e.g. the name of a celebrity; contrast this to unsafe content, which is a more abstract concept). Applying LoRA fine-tuning after UCE unlearning, we find no evidence of concept resurgence. We then instead fine-tune the full set of parameters, which yields a small degree of resurgence. Finally, motivated by this result, we choose to fully fine-tune the cross-attention layers only. We see that the resurgence is comparable between the two (Table˜3), suggesting that the nature of UCE’s closed-form edit being localized to the cross-attention layers may make it very robust.
@@
The second difference between the seven algorithms is the subset of model parameters that are updated in the unlearning step. Section˜3 focuses primarily on modifying the either the cross-attention layers (with the exception of MACE, which also updates the rest of the model parameters via LoRA fine-tuning) or the automatically selected parameter subset (i.e. SalUn and SHS). Here, we focus on ESD in the single celebrity erasure task and the copyright erasure task, which both exhibit very high degrees of concept resurgence. In each of these tasks, we vary the subset of parameters that are updated in the unlearning step: either all of the parameters, all of the parameters except those in the cross-attention layers, and only those in the cross-attention layers. We find that the cross-attention parameters do indeed play the most important role in unlearning for these tasks and that unlearning on all the parameters only provided marginal gains in preventing resurgence (Fig.˜16).
Finetuning Hyperparameters Finally, we investigate how hyperparameter choices such as dataset size and number of fine-tuning steps impact the severity of resurgence. In Appendix˜J, we show that even with much smaller amounts of data or smaller amounts of fine-tuning steps that resurgence still occurs. For example, for MACE on our erase 10 celebrities task, only 20 fine-tuning steps are necessary for resurgence to occur with 250 samples.
@@ 5 Why Does Concept Resurgence Occur?
To better understand the conditions under which forgotten concepts can resurface during fine-tuning, we analyze a linear score-based diffusion model. Our results illustrate that any non-zero overlap between the subspace of the forgotten concept and the subpsace spanned by the fine-tuning gradients is sufficent to induce resurgence in this setting. Let $\mathcal{C}\subset\mathbb{R}^{d}$ denote the erased concept subspace and let $\mathcal{D}_{\rm FT}$ be the fine-tuning dataset whose per-example gradients span a subspace $\mathcal{S}$. We assume that the orthogonal projection of $\mathcal{S}$ onto $\mathcal{C}$, denoted $P_{\mathcal{C}}(\mathcal{S})$, is nonzero (i.e., the subspaces are not orthogonal). This assumption reflects the realistic likelihood of incidental alignment in high-dimensional settings. Although this overlap is sufficient to trigger some degree of resurgence, it does not account for the magnitude of the effect. Our goal is to characterize how this seemingly weak interaction can be amplified into meaningful resurgence. In particular, our analysis identifies two bounds that characterize how resurgence occurs in this model:
  • Gradient resurgence bound. This bound identifies when fine-tuning gradients reappear in the forgotten subspace $\mathcal{C}$, despite prior unlearning. It shows that nonzero gradient mass arises in $\mathcal{C}$ whenever there is residual alignment between the fine-tuning subspace $\mathcal{S}$ and $\mathcal{C}$. Formally:
    $\left\|P_{\mathcal{C}}\left(\nabla_{W}\mathcal{L}_{t}\right)\right\|_{F}\geq 2\sqrt{1-\alpha_{t}}\cdot\sqrt{\gamma(\mathcal{S},\mathcal{C})},$
    where $\gamma(\mathcal{S},\mathcal{C})\triangleq\lambda_{\min}(P_{\mathcal{S}}P_{\mathcal{C}}P_{\mathcal{S}})$ measures the worst-case leakage from $\mathcal{S}$ into $\mathcal{C}$. This overlap ensures that even when concepts in $\mathcal{C}$ have been suppressed, fine-tuning gradients computed from noise-perturbed data can reintroduce them if they are not fully orthogonal to the directions encoded in the new task. Notably, this bound is most active at early diffusion timesteps, where $1-\alpha_{t}$ is large and thus amplifies the residual error when there is any amount of overlap.
  • Curvature-limited sensitivity. This bound captures the model’s geometric sensitivity to reactivation. Even if gradient mass in $\mathcal{C}$ is small, the induced update can be large if the curvature in those directions is low. Formally, for any update $\Delta W$ supported in $\mathcal{C}$, we have:
    $\|P_{\mathcal{C}}\Delta W\|_{F}\geq\frac{2\sqrt{1-\alpha_{t}}\cdot\sqrt{\gamma(\mathcal{S},\mathcal{C})}}{\alpha_{t}\lambda_{\max}^{{\mathcal{C}}}+(1-\alpha_{t})},$
    where $\lambda_{\max}^{{\mathcal{C}}}:=\lambda_{\max}(P_{\mathcal{C}}\Sigma P_{\mathcal{C}})$ is the maximum variance in the forgotten subspace. This bound reveals a key amplification mechanism: low-variance directions are highly sensitive to reactivation, since small gradients can produce large updates when curvature is shallow.
Proposition 5.1 (Linear diffusion model resurgence).
Assume a linear diffusion model with residual of the form
$\epsilon_{W}(x_{t},t):=Wx_{t}-\epsilon$
for some matrix $W\in\mathbb{R}^{d\times d}$, where $\epsilon\sim\mathcal{N}(0,I)$ is independent Gaussian noise. Let $\mathcal{C}\subset\mathbb{R}^{d}$ be a subspace, and let $\mathcal{D}_{\mathrm{FT}}$ be a fine-tuning dataset whose induced gradient directions span a subspace $\mathcal{S}$. Let $P_{\mathcal{C}}=U_{\mathcal{C}}U_{\mathcal{C}}^{\top}$ and $P_{\mathcal{S}}=U_{\mathcal{S}}U_{\mathcal{S}}^{\top}$ denote the orthogonal projection matrices onto $\mathcal{C}$ and $\mathcal{S}$, respectively. Define the leakage
$\gamma(\mathcal{S},{\mathcal{C}})\triangleq\lambda_{\min}(P_{\mathcal{S}}P_{\mathcal{C}}P_{\mathcal{S}}).$
Let $x_{0}\sim\mathcal{D}_{\mathrm{FT}}$ have covariance $\Sigma$, and define the forward-corrupted input as
$x_{t}:=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon\quad\text{so that}\quad\Sigma_{t}:=\mathbb{E}[x_{t}x_{t}^{\top}]=\alpha_{t}\Sigma+(1-\alpha_{t})I.$
Let $\lambda_{\max}^{\mathcal{C}}\triangleq\lambda_{\max}(P_{\mathcal{C}}\Sigma P_{\mathcal{C}})$ and suppose $P_{\mathcal{C}}\mathcal{S}\neq 0$. Then if the prior unlearning was successful, i.e. $P_{\mathcal{C}}W=0$, we obtain bounds characterizing resurgence:
  1. 1.
    Gradient resurgence: The fine-tuning gradient projected into $\mathcal{C}$ satisfies:
    $\left\|P_{\mathcal{C}}\left(\nabla_{W}\mathcal{L}_{t}\right)\right\|_{F}\geq 2\sqrt{1-\alpha_{t}}\cdot\sqrt{\gamma(\mathcal{S},\mathcal{C})}.$
  2. 2.
    Curvature-limited sensitivity: The update $\Delta W\in\mathbb{R}^{d\times d}$ in the weight matrix supported in the forgotten subspace $\mathcal{C}$ satisfies:
    $\|P_{\mathcal{C}}\Delta W\|_{F}\geq\frac{2\sqrt{1-\alpha_{t}}\cdot\sqrt{\gamma(\mathcal{S},\mathcal{C})}}{\alpha_{t}\lambda_{\max}^{\mathcal{C}}+(1-\alpha_{t})}.$
We provide a proof of the gradient resurgence bound in Appendix G and the proof of the curvature-limited sensitivity bound to Appendix H. The basic idea for gradient bound is to notice the norm of the gradient is lower bounded by the norm of the Frobenius norm $A:=\mathbb{E}[\epsilon_{W}(x_{t},t)x_{t}^{\top}]$ restricted to $\mathcal{C}$ multiplied by the overlap term $\gamma(\mathcal{S},\mathcal{C})$ which simplifies to $1-\alpha_{t}$ when excact unlearning has occurred. The bound on the update uses a bound that relies on the loss function for the weight $W$ being quadratic.
Together, these bounds clarify the structural and dynamical factors that govern resurgence after unlearning. Specifically, we identify two distinct contributors: (1) conceptual overlap, captured by the gradient resurgence bound, quantifies when fine-tuning gradients reappear in the forgotten subspace $\mathcal{C}$. The bound depends on the overlap between $\mathcal{C}$ and the supervision subspace $\mathcal{S}$, via the leakage term $\gamma(\mathcal{S},\mathcal{C})$, and scales linearly with $\sqrt{1-\alpha_{t}}$, reflecting the increased influence of noise-perturbed input at early diffusion steps. In the simplified setting where $P_{\mathcal{C}}W=0$, this term isolates reactivation due purely to residual alignment. More generally, when signal is injected into $W$, the bound acquires an additional contribution proportional to $P_{\mathcal{C}}W\Sigma$, which can dominate in late timesteps and further amplify resurgence; (2) amplification, captured by the curvature-limited sensitivity bound, governs how strongly the model responds to gradient mass in $\mathcal{C}$. Even small gradients can induce large parameter updates when the loss curvature is low. This effect is most pronounced when $\mathcal{C}$ aligns with low-variance directions in the data (i.e., small $\lambda_{\max}^{\mathcal{C}}$), and when $\alpha_{t}$ is large, but not too close to 1 so that the noise has diminished but curvature remains anisotropic. In contrast, early timesteps ($\alpha_{t}\ll 1$) introduce strong isotropic curvature, suppressing updates and making reactivation less likely.
@@ 6 Discussion and Limitations
Our investigation opens several important directions for future work. First, our theoretical analysis is restricted to the linear setting, and it remains an open question whether similar characterizations of concept resurgence extend to nonlinear models. Exploring such extensions could inform new strategies for mitigating resurgence and improving the robustness of unlearning procedures. Second, our empirical evaluation is limited to standard academic benchmarks and synthetic settings. Further research is needed to assess the practical relevance of concept resurgence in real-world deployments, particularly in scenarios involving long-horizon or compositional fine-tuning, where interleaved updates may amplify vulnerabilities.
Concept resurgence also raises important questions about responsibility for downstream harms. Despite a developer’s best efforts to sanitize a model using these techniques, a downstream user who fine-tunes a published model might be surprised to discover that guardrails put in place by the developer no longer exist. This creates a dilemma: is the developer obligated to permanently and irrevocably erase problematic concepts, or does responsibility shift to the downstream if they (inadvertently) reintroduce them? Despite these challenges, concept unlearning remains a valuable tool for model developers. By identifying its vulnerabilities, our work aims to drive the development of erasure techniques that remain robust throughout a model’s life-cycle or develop tools that can help developers anticipate when concept resurgence is likely to happen.
@@ Appendix A Additional Qualitative Examples
In this section, we include qualitative results for the copyright and object erasure tasks in Figure˜5 and Figure˜6, respectively. These results are analogous to those presented in Figure˜2 for the celebrity erasure task. We choose to exclude qualitative examples of resurgence for the unsafe content task, as these may be upsetting. For a quantitative evaluation of this task across all seven unlearning algorithms, we refer readers to Table˜1.
@@
Figure 7: Selected images generated by SD v1.4 after initially applying each new unlearning algorithm (top row) and after subsequent fine-tuning (bottom row) in the celebrity unlearning task. In each case, the model initially unlearns the target concept; e.g., how to generate images of Andrew Garfield. However, fine-tuning on unrelated images still inadvertently reintroduces the erased concepts on these new baselines.
Figure 8: Selected images generated by SD v1.4 after initially applying each new unlearning algorithm (top row) and after subsequent fine-tuning (bottom row) in the copyright unlearning task. In each case, the model initially unlearns the target concept; e.g., how to generate images of Iron Man. However, fine-tuning on unrelated images still inadvertently reintroduces the erased concepts on these new baselines.
Figure 9: Selected images generated by SD v1.4 after initially applying each new unlearning algorithm (top row) and after subsequent fine-tuning (bottom row) in the object unlearning task. In each case, the model initially unlearns the target concept; e.g., how to generate images of an automobile. However, fine-tuning on unrelated images still inadvertently reintroduces the erased concepts on these new baselines.